OpenStudy (anonymous):

Which of the following is a step in simplifying the expression x multiplied by y to the power of 4 over x to the power of negative 5 multiplied by y to the power of 5, the whole to the power of negative 3.? x to the power of negative 3 multiplied by y, the whole over x to the power of negative 8 multiplied by y to the power of 2. x to the power of negative 3 multiplied by y to the power of negative 12, the whole over x to the power of negative 5 multiplied by y to the power of 5. x to the power of negative 3 multiplied by y, the whole over x to the power of negative 5 multiplied by y to the p

OpenStudy (anonymous):

@uChezzy

OpenStudy (anonymous):

@Umangiasd

OpenStudy (anonymous):

Okay, to start with, we'll write it in a more math way xd \[(\dfrac {(x \times y)^4} {x^5 \times y^5} ) ^{-3}\] Now, to simplify this i would recal the property used in the last problem but in the reverse way, so you'll have \[(\dfrac {(x \times y)^4} {(x \times y)^5} ) ^{-3}\] This expression can be treated with the division power property (don't know the proper english name) \[\dfrac { a^n} { a^m} = a^{n-m}\] Let me see what you get c: